a large shipment of items contains 8% defective items. four items of the shipment are randomly selection. what is the chance that the selection will include at least one defective item?
Answer by stanbon(75887)

a large shipment of items contains 8% defective items. four items of the shipment are randomly selection. what is the chance that the selection will include at least one defective item?
---
Binomial Problem with n = 4 and p(defective) =0.08 ; p(not defective) = 0.92
--------------
P(1<= x <=4) = 1 - P(x = 0)
= 1 - (0.92)^4
= 1 - 0.7164
= 0.2836
=============
Cheers,
Stan H.
=============